OpenStudy (anonymous):

Will Medal!!! The length of a rectangle is 2 more than 5 times its width. The perimeter of the rectangle is 88 inches. What is the length of the rectangle? 32 in. 37 in. 42 in. 47 in.

OpenStudy (superflytyguy):

The perimeter of a rectangle is 2L + 2W. Substitute in your other variables. 2(5 + 2x) + 2(x) = perimeter 10 + 4x + 2x = 88 10 + 6x = 88 6x = 78 x = 13 We stated earlier that x = the width. Therefore, the width is 13 feet. Length = 5 + 2x = 5 + 2(13) = 5 + 26 = 31

OpenStudy (superflytyguy):

Where 32 come from... idk

OpenStudy (catlover5925):

ok so I got 37...I tried them all but with 37 I added 37 + 37 since there are 2 sides, then I got 74 after that I took 88(whole perimeter) and subtracted it from 74 and I got14 which is both the sides together so I divided that by 2 and got 7 if you take 7*5+2 you get 37 which is right

OpenStudy (catlover5925):

@SuperflyTyGuy and @maybubble does that make sense to both of you guys

OpenStudy (superflytyguy):

You're right @Catlover5925